A collection of 1 cent (pennies), 5 cent (pickels) and 10 cent (dimes) totals 2.12.00. The total number of coins is 42 and there are twice as many 5 cent (nickels0 as 10 cent (dimes). How many of each kind are there?
Answer by KnightOwlTutor(293)

X=number of dimes
2X=number of nickels
42-x-2x=number of pennies
.10x+.05(2x)+.01(42-3x)=2.12
To make this expression easier to deal with multiply each term by 100
10x+10x+42-3x=212
Simplify by combining like terms
(10+10-3)x+42=212
17x+42=212
Subract 42 from both sides
17X=170
X=10=number of dimes
2X=20=number of nickels
42-3x=12=number of pennies


Let's check our answer 10(.1)=$1.00
20(.05)=$1.00
12(.01)=$0.12
add together and come up with $2.12